# Ancestor regression in structural vector autoregressive models

### v1

@@ 1 Introduction
Real-world datasets often exhibit a time structure violating the i.i.d. assumption widely used in causal discovery and beyond. Such data can be modeled with (structural) vector autoregressive models, i.e., using past and current observations of the time series as predictors. While the time dependence implies certain difficulties in estimation, it offers some advantages as well because a predictor cannot causally affect other variables that represent earlier time points. With independent innovation terms, identifiability guarantees as for fully independent observations can be found under similar structural assumptions, see Peters et al., (2013).
@@ 1.2 Structural vector autoregressive model
Let the $\tau$-lagged causal ancestors of $x_{t,j}$, $\text{AN}^{\tau}\mathopen{}\mathclose{{}\left(j}\right)$ be all $k$ for which there exists a directed path from $x_{t-\tau,k}$ to $x_{t,j}$ in the full causal graph. Analogously, we say $k\in\text{PA}^{\tau}\mathopen{}\mathclose{{}\left(j}\right)\iff j\in\text{CH}^{\tau}\mathopen{}\mathclose{{}\left(k}\right)$ to denote parents and children if there is an edge from $x_{t-\tau,k}$ to $x_{t,j}$. For $\tau=0$, $\text{AN}^{\tau=0}(j)$ are the instantaneous ancestors of $x_{t,j}$.
@@ 1.3 Identifying $AN^{\tau}$ via ancestor regression
The resulting least squares coefficients are (in population) 00 for all non-ancestors, while they are - up to few counter-examples - non-zero for the ancestors. Hence, one can identify the ancestors using a simple least squares regression. We will show here how this method can be extended to the related SVAR (1).
@@
There can be cases where $\beta_{k}^{f,j,\tau}=0$ although $k\in\text{AN}^{\tau}\mathopen{}\mathclose{{}\left(j}\right)$. For some data generating mechanisms, this happens regardless of the choice of function $f\mathopen{}\mathclose{{}\left(\cdot}\right)$. We want to characterize these cases. Denote the Markov boundary of $x_{t,k}$ by $\text{MA}\mathopen{}\mathclose{{}\left(k}\right)$ and get the corresponding vector as
@@
i.e., the instantaneous children through which a directed path goes to $x_{t+\tau,j}$. Then, we call $\text{MA}^{\tau\rightarrow j}\mathopen{}\mathclose{{}\left(k}\right)$ the restricted Markov boundary and get the corresponding vector as
@@
The implied conditional independence in Theorem 2 means that there are no directed paths from $x_{t,k}$ to $x_{t+\tau,j}$ that do not go through some other $x_{t,l}$, or that all these paths cancel each other out. It is trivially always fulfilled for instantaneous effects, i.e., for these, the result is very similar to Theorem 3 of Schultheiss and Bühlmann, (2023) and the same corresponding examples apply, i.e., Gaussian error terms and identical contribution between predictor and noise, see there for details. Accordingly, if all paths from a lagged ancestor start with an undetectable immediate effect, this lagged effect cannot be detected either.
@@ 2 Estimation from data and asymptotics
$\hat{\boldsymbol{\xi}}^{\tau}$ $=\begin{pmatrix}\hat{\boldsymbol{\xi}}^{\tau}_{1}&\ldots&\hat{\boldsymbol{\xi}}^{\tau}_{p}\end{pmatrix}^{\top},$
@@
where $\Phi\mathopen{}\mathclose{{}\left(\cdot}\right)$ denotes the cumulative distribution function of the standard normal distribution.
@@ 2.1 Simulation example
On the left-hand side of Figure 1, we see the average absolute z-statistics for ancestors and non-ancestors for the different sample sizes. We distinguish between three types of ancestors: instantaneous ancestors, lagged ancestors for which $\tilde{\mathbf{B}}_{4,k}\neq 0$, and lagged ancestors from which all causal paths start with an instantaneous effect. The last are the hardest to detect. Otherwise, lagged ancestors have stronger signals than instantaneous ones. This agrees with the intuition that it is easier to find a directed causal path if it is a priori known that only one direction could be possible. For non-ancestors, the observed average of the absolute z-statistics is close to the theoretical mean under the asymptotic null distribution as desired. On the right-hand side, we see that we can control the type I error at the desired level for every sample size. As expected, the power to detect ancestors increases with larger sample sizes. However, driven by this last group of ancestors, there are still some undetected ancestors for $T=10^{6}$. For the other groups, we obtain almost perfect power. One could then also infer these missed effects by recursive arguments. We discuss this for the case of networks below.
@@ 3.1 Instantaneous effects
If we make type I errors, this could create contradictions leading to cycles. Then, we gradually decrease the significance level for edges within these cycles until no more cycles remain. The largest significance level for which no loops occur is also an asymptotically valid p-value for the null hypothesis that the data come from model (2) with assumptions (A1) - (A4) as we have asymptotic error control under this model class. An example where $\alpha=0.05$ leads to cycles is demonstrated in Figure 2. We keep the edge from $x_{t,2}$ to $x_{t,4}$ using the initial significance level as it is not part of any cycles.
@@ 3.2 Summary time graph
To obtain it, we first assign a p-value to each potential edge $k\rightarrow j$, say, $p_{k}^{j}$. There are $p+1$ p-values corresponding to this edge, i.e., $p_{k}^{j,0},\ldots p_{k}^{j,p}$. We combine them using ideas from Meinshausen et al., (2009) designed to combine p-values under arbitrary dependence. For details, see Appendix A.7. Again, we apply the Bonferroni-Holm multiplicity correction to control the family-wise error rate. This allows to add recursively more edges while still controlling the type I error rate. Model (1) does not imply that the summary graph must be acyclic. Thus, we output the result after the recursive construction even in the presence of cycles. For example, in Figure 2 if the depicted p-values are (multiplicity corrected) summary p-values, the obtained summary graph is the second to the left.
@@ 3.3 Simulation example
Figure 3: Nodewise ancestor detection in a structural vector autoregressive model of order $p=1$ with $6$ variables. The results are based on $1000$ simulation runs. Depicted is the family-wise error rate of false causal detection versus the fraction of detected ancestors. The curves use the level of the test $\alpha$ as implicit curve parameter. The pluses correspond to nominal $\alpha=5\%$. The vertical line is at actual $5\%$. We consider the different sample sizes $10^{2}$ (solid, black), $10^{3}$ (dashed, red), $10^{4}$ (dotted, green), $10^{5}$ (dot-dashed, blue), and $10^{6}$ (long-dashed, pink). On the left: instantaneous effects. On the right: summary graph.
@@ 5.2 Conclusion
We also obtain asymptotic power up to few pathological cases. In networks, additional effects can be inferred by the logic that an ancestor of an ancestor must be an ancestor. In our simulation, we see that this can help to find almost all ancestors without errors even when some connections are individually hard to find at a given sample size.

### v2

@@ 1 Introduction
Real-world datasets often exhibit a time structure violating the i.i.d. assumption widely used in causal discovery and beyond. Such data can be modelled with (structural) vector autoregressive models, i.e., using past and current observations of the time series as predictors. While the time dependence implies certain difficulties in estimation, it also offers some advantages because a predictor cannot causally affect other variables that represent earlier time points. With independent innovation terms, identifiability guarantees as for fully independent observations can be found under similar structural assumptions, see Peters et al., (2013).
@@ 1.2 Structural vector autoregressive model
Instantaneous effects are often assumed to represent processes that happen at a higher frequency than the observation sampling frequency. As such it is not clear that they must be acyclic. However, the assumption is standard when extending causal discovery algorithms to applications on time series data, see, e.g., Peters et al., (2013). We follow this proposal and take the assumption as a given. However, as cyclic effects from high-frequency processes can be attributed to unobserved variables, the impact is implicitly covered by our simulation study in Section 3.3.
@@
Let the $\tau$-lagged causal ancestors of $j$, $\text{AN}^{\tau}\mathopen{}\mathclose{{}\left(j}\right)$ be all $k$ for which there exists a directed path from $x_{t-\tau,k}$ to $x_{t,j}$ for all $t$ in the full causal graph. Analogously, we say $k\in\text{PA}^{\tau}\mathopen{}\mathclose{{}\left(j}\right)\iff j\in\text{CH}^{\tau}\mathopen{}\mathclose{{}\left(k}\right)$ to denote parents and children if there is an edge from $x_{t-\tau,k}$ to $x_{t,j}$ for all $t$. For $\tau=0$, $\text{AN}^{\tau=0}(j)$ are the instantaneous ancestors of $j$. As the $\mathbf{B}_{\tau}$ are time-invariant, we also omit the time index for ancestors, children, and parents.
@@ 1.3 Identifying $AN^{\tau}$ via ancestor regression
The resulting least squares coefficients are (in population) 00 for all non-ancestors, while they are - up to few counter-examples - non-zero for the ancestors. For a discussion of such counter-examples, see Section 2.2 in Schultheiss and Bühlmann, (2023), we adapt this to time series in Theorem 2. The most important exception is Gaussian noise. If there are adjacent variables with Gaussian noise, several causal orders could imply the same multivariate distribution. Due to the type I guarantees, ancestor regression cannot detect an ancestral relation in such cases and remains conservative. Accordingly, the related LiNGAM algorithm (Shimizu et al.,, 2006) allows for at most one Gaussian noise term by assumption - otherwise it becomes inconsistent.
Based on the observation above, one can identify the ancestors, or at least some of them, using a simple least squares regression. We will show here how this method can be extended to the related SVAR (1).
@@
There can be cases where $\beta_{k}^{f,j,\tau}=0$ although $k\in\text{AN}^{\tau}\mathopen{}\mathclose{{}\left(j}\right)$. For some data generating mechanisms, this happens regardless of the choice of function $f\mathopen{}\mathclose{{}\left(\cdot}\right)$. We want to characterize these cases. For instantaneous effects, these exceptions are - non-surprisingly - in close correspondence to the i.i.d.  case. For lagged effects, for which the causal paths do not start with an instantaneous effect, it is related to faithfulness (Spirtes et al.,, 2000, Chapter 2.3.3). If the target is independent of its lagged ancestor due to the cancellation of terms, we cannot detect this relationship.
For the precise statements, we introduce some notation. Denote the Markov boundary of $x_{t,k}$ by $\text{MA}\mathopen{}\mathclose{{}\left(k}\right)$ and get the corresponding vector as
@@
i.e., if $l\in\text{CH}^{0,\tau\rightarrow j}\mathopen{}\mathclose{{}\left(k}\right)$, there is a path $x_{t,k}\rightarrow x_{t,l}\rightarrow\ldots\rightarrow x_{t+\tau,j}$ for all $t$. Then, we call $\text{MA}^{\tau\rightarrow j}\mathopen{}\mathclose{{}\left(k}\right)$ the restricted Markov boundary and get the corresponding vector as
@@
We provide some intuition for the necessary conditional independence. It is always fulfilled for instantaneous effects as the restricted Markov boundary blocks all mediating or confounding paths. Hence, the sufficient condition boils down to the first equality only which is in correspondence to Theorem 3 of Schultheiss and Bühlmann, (2023) and the same corresponding examples apply, i.e., Gaussian error terms and identical contribution between predictor and noise, see there for details.
For a lagged effect, if there are paths from $x_{t,k}$ to $x_{t+\tau,j}$ that do not go through some other $x_{t,l}$, conditional independence only holds if those paths cancel each other out, i.e., the data is unfaithful. If there are only such paths, $\text{CH}^{0,\tau\rightarrow j}\mathopen{}\mathclose{{}\left(k}\right)=\emptyset$. This implies that $\text{MA}^{\tau\rightarrow j}\mathopen{}\mathclose{{}\left(k}\right)$ contains only ancestors such that the conditional expectation is indeed linear. Thus, in that case, unfaithfulness becomes sufficient and necessary.
Paths that go through some other $x_{t,l}$, i.e., starting with an immediate effect are a composition of the two cases above. The effects are detectable if both, the effect from $x_{t,k}$ to $x_{t,l}$ and from $x_{t,l}$ to $x_{t+\tau,k}$ are as well.
@@ 2 Estimation from data and asymptotics
$\hat{\boldsymbol{\xi}}^{\tau}$ $=\begin{pmatrix}\hat{\boldsymbol{\xi}}^{\tau}_{1}&\ldots&\hat{\boldsymbol{\xi}}^{\tau}_{d}\end{pmatrix}^{\top},$
@@
where $\Phi\mathopen{}\mathclose{{}\left(\cdot}\right)$ denotes the cumulative distribution function of the standard normal distribution. We summarize the procedure in Algorithm 1. Note that we sometimes suppress explicit time indexing for easier readability.
@@
Algorithm 1 Ancestor regression in SVAR
Inputs:
      $\mathbf{x}\in\mathbb{R}^{T\times d}$, degree $p$, target $j$, nonlinearity $f:\mathbb{R}\to\mathbb{R}$
$\hat{\boldsymbol{\xi}}$leftarrow project out the contribution of $\mathbf{x}_{t-1}$ to $\mathbf{x}_{t-p}$ from $\mathbf{x}_{t}$
for $\tau=0$ to $p$ do
     $\hat{\boldsymbol{\xi}}^{\tau}$leftarrow project out the contribution of $\mathbf{x}_{t-\tau-1}$ to $\mathbf{x}_{t-\tau-p}$ from $\mathbf{x}_{t}$
     $\mathbf{s}^{j,\tau}\leftarrow z$ values from OLS$\mathopen{}\mathclose{{}\left(f\mathopen{}\mathclose{{}\left(\hat{\boldsymbol{\xi}}_{j}^{\tau}}\right)\sim\hat{\boldsymbol{\xi}}}\right)$, where $\hat{\boldsymbol{\xi}}$ is appropriately lagged by $\tau$ steps
     for $k=1$ to $d$ do
         $p_{k}^{j,\tau}=2\mathopen{}\mathclose{{}\left\{1-\Phi\mathopen{}\mathclose{{}\left(|s_{k}^{j,\tau}|}\right)}\right\}$
Returns:
      $p_{k}^{j,\tau}\ \forall k,\tau$
@@ 2.1 Simulation example
On the left-hand side of Figure 1, we see the average absolute z-statistics for ancestors and non-ancestors for the different sample sizes. We distinguish between three types of ancestors: instantaneous ancestors, lagged ancestors for which $\tilde{\mathbf{B}}_{4,k}\neq 0$, and lagged ancestors from which all causal paths start with an instantaneous effect. Figure 2 illustrates the different types of lagged ancestors. The path from $x_{t-1,4}$ to $x_{t,4}$ is not mediated by instantaneous effects at $t-1$ and $\mathopen{}\mathclose{{}\left(\tilde{\mathbf{B}}_{1}}\right)_{4,4}\neq 0$. Such lagged ancestors have stronger signals than instantaneous ones. This agrees with the intuition that it is easier to find a directed causal path if it is a priori known that only one direction could be possible, i.e., the lagged variable can be an ancestor but not a descendant such that one type of dependence is excluded. However, the only causal path from $x_{t-1,3}$ to $x_{t,4}$ starts with an instantaneous effect such that $\mathopen{}\mathclose{{}\left(\tilde{\mathbf{B}}_{1}}\right)_{4,3}=0$. To detect such an effect, one must essentially detect that instantaneous effect but with a target that contains additional noise from all other contributions. Hence, this is harder than detecting instantaneous effects. As we randomly sample the edge weights from a continuous distribution, there are no unfaithful zeros in $\tilde{\mathbf{B}}$.
$x_{t-1,3}$$x_{t-1,4}$$x_{t,3}$$x_{t,4}$
Figure 2: Example of the different types of ancestors in Figure 1.
For non-ancestors, the observed average of the absolute z-statistics is close to the theoretical mean under the asymptotic null distribution as desired. On the right-hand side, we see that we can control the type I error at the desired level for every sample size. As expected, the power to detect ancestors increases with larger sample sizes. However, driven by the last group of ancestors discussed above, there are still some undetected ancestors for $T=10^{6}$. For the other groups, we obtain almost perfect power. One could then also infer these missed effects by recursive arguments. We discuss this for the case of networks below.
@@ 3.1 Instantaneous effects
Algorithm 2 Instantaneous graph
Inputs:
      $\mathbf{x}\in\mathbb{R}^{T\times d}$, degree $p$, nonlinearity $f:\mathbb{R}\to\mathbb{R}$, level $\alpha$
for $j=1$ to $d$ do
     Use Algorithm 1 with target $j$ to get $p_{k}^{j,0}\ \forall k\neq j$
Apply Bonferroni-Holm correction over all $d\mathopen{}\mathclose{{}\left(d-1}\right)$ tests to obtain multiplicity corrected $P_{k}^{j,0}$
Initiate $\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right)=\emptyset\forall j$
for $j,k=0$ to $d$ do
     if $P_{k}^{j,0}<\alpha$ then
         $\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right)=\mathopen{}\mathclose{{}\left\{\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right),k}\right\}$triangleright Find initial estimated set using significant tests
while $\exists j,k,l$ such that $k\in\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right)$, $l\in\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(k}\right)$ but $l\notin\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right)$ do
     $\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right)=\mathopen{}\mathclose{{}\left\{\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right),l}\right\}$triangleright Add ancestors recursively
Initiate $\tilde{\alpha}=\alpha$
while $\exists j$ such that $j\in\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right)$ do
     $I=\mathopen{}\mathclose{{}\left\{j:\quad j\in\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right)}\right\}$triangleright Variables that are on a cycle
     Find highest p-value $\tilde{\alpha}\coloneqq P_{l}^{k,0}<\alpha$, where $k,l\in I$
     Re-estimate ancestor relations within I using level $\tilde{\alpha}$
Returns:
      $\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right)\quad\forall j$ and $\tilde{\alpha}$
@@
If we make type I errors, this could create contradictions leading to cycles. Then, we gradually decrease the significance level for edges within these cycles until no more cycles remain. The largest significance level for which no loops occur is also an asymptotically valid p-value for the null hypothesis that the data come from model (2) with assumptions (A1) - (A4) as we have asymptotic error control under this model class. An example where $\alpha=0.05$ leads to cycles is demonstrated in Figure 3. We keep the edge from $x_{t,2}$ to $x_{t,4}$ using the initial significance level as it is not part of any cycles. The procedure is summarized in Algorithm 2.
@@ 3.2 Summary time graph
Algorithm 3 Summary graph
Inputs:
      $\mathbf{x}\in\mathbb{R}^{T\times d}$, degree $p$, nonlinearity $f:\mathbb{R}\to\mathbb{R}$, level $\alpha$
for $j=1$ to $d$ do
     Use Algorithm 1 with target $j$ to get $p_{k}^{j,\tau}\ \forall k\neq j,\tau\leq p$
     for $k\neq j$ do
         Obtain $p_{k}^{j}$ from $p_{k}^{j,0},\ldots,p_{k}^{j,p}$ as in Appendix A.7
Apply Bonferroni-Holm correction over all $d\mathopen{}\mathclose{{}\left(d-1}\right)$ tests to obtain multiplicity corrected $P_{k}^{j}$
Initiate $\widehat{\text{AN}}\mathopen{}\mathclose{{}\left(j}\right)=\emptyset\forall j$
for $j,k=0$ to $d$ do
     if $P_{k}^{j,0}<\alpha$ then
         $\widehat{\text{AN}}\mathopen{}\mathclose{{}\left(j}\right)=\mathopen{}\mathclose{{}\left\{\widehat{\text{AN}}\mathopen{}\mathclose{{}\left(j}\right),k}\right\}$triangleright Find initial estimated set using significant tests
while $\exists j,k,l$ such that $k\in\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(j}\right)$, $l\in\widehat{\text{AN}}^{0}\mathopen{}\mathclose{{}\left(k}\right)$ but $l\notin\widehat{\text{AN}}\mathopen{}\mathclose{{}\left(j}\right)$ do
     $\widehat{\text{AN}}\mathopen{}\mathclose{{}\left(j}\right)=\mathopen{}\mathclose{{}\left\{\widehat{\text{AN}}\mathopen{}\mathclose{{}\left(j}\right),l}\right\}$triangleright Add ancestors recursively
Returns:
      $\widehat{\text{AN}}\mathopen{}\mathclose{{}\left(j}\right)\quad\forall j$
@@
To obtain it, we first assign a p-value to each potential edge $k\rightarrow j$, say, $p_{k}^{j}$. There are $p+1$ p-values corresponding to this edge, i.e., $p_{k}^{j,0},\ldots p_{k}^{j,p}$. We combine them using ideas from Meinshausen et al., (2009) designed to combine p-values under arbitrary dependence. For details, see Appendix A.7. Again, we apply the Bonferroni-Holm multiplicity correction to control the family-wise error rate. This allows to add recursively more edges while still controlling the type I error rate. Model (1) does not imply that the summary graph must be acyclic. Thus, we output the result after the recursive construction even in the presence of cycles. For example, in Figure 3 if the depicted p-values are (multiplicity corrected) summary p-values, the obtained summary graph is the second to the left. We summarize the procedure in Algorithm 3.
@@ 3.3 Simulation example
Figure 4: Nodewise ancestor detection in a structural vector autoregressive model of order $p=1$ with $6$ variables. The results are based on $1000$ simulation runs. Depicted is the family-wise error rate of false causal detection versus the fraction of detected ancestors. The curves use the level of the test $\alpha$ as an implicit curve parameter. The pluses correspond to nominal $\alpha=5\%$. The vertical line is at actual $5\%$. The other symbols represent the performance of the LiNGAM algorithm. On the left: instantaneous effects. On the right: summary graph.
@@
Additionally, we show on the left side the performance of the LiNGAM algorithm as described in Hyvärinen et al., (2010). For this, we use the code published together with Moneta et al., (2013). As the LiNGAM algorithm by default does not search for sparse estimates $\hat{\mathbf{B}}_{\tau}$ for $\tau>0$, a straightforward comparison is only possible for instantaneous effects. We note that the LiNGAM algorithm leads to higher power, especially for low sample sizes. However, it does not allow for an interpretable error control at a predefined level. Similar results were obtained for the comparison in the i.i.d. case, see Schultheiss and Bühlmann, (2023).
In the next experiment, we analyze the effect of unobserved variables. We reuse the same simulation setting but randomly consider one of the time series as unobserved for each simulation run. We run ancestor regression and LiNGAM with the observed time series and calculate the same performance statistics. The results are in Figure 5.
Control of the familywise error rate at a fixed level no longer works for ancestor regression. However, there is still a separation between ancestors and non-ancestors in terms of the effect size. For long time series, there is a sweet spot with high power at a fairly low error rate. Hence, the ordering of the p-values still gives some indication of what could be the true ancestors. Some of the curves do not start at 00 power and error rate as there are p-values that are numerically equal to 00 such that no reasonable ordering can be made for these observations.
At our target level $\alpha=0.05$, the power remains comparable to the case without hidden variables or is even increased for some sample sizes. The unobserved variables are mainly a problem for error control as the assumptions are not fulfilled but not for detection per se. Similarly, the power of the LiNGAM algorithm remains high while the error rate is increased.
Additional simulation results with higher numbers of time series, both with and without unobserved variables are in Appendix B.2.
Figure 5: Nodewise ancestor detection in a structural vector autoregressive model of order $p=1$ with $6$ variables, whereof one is unobserved. See the caption of Figure 4 for explanations.
@@ 5.2 Conclusion
We also obtain asymptotic power up to few pathological cases. In networks, additional effects can be inferred by the logic that an ancestor of an ancestor must be an ancestor. In our simulation, we see that this can help to find almost all ancestors without errors even when some connections are individually hard to find at a given sample size. When incorporating unobserved time series, which is a violation of our assumptions, error control at a predefined level does not work anymore. However, the ordering of the test statistics still provides some indication of what could be true ancestors.
@@
Code scripts to reproduce the results presented in this paper are available here
https://github.com/cschultheiss/SVAR-Ancestor.